OpenStudy (anonymous):

How do I solve by factoring: 4x^2 = 9

hero (hero):

4x^2 = 9 x^2 = 9/4 x = sqrt(9/4) x = 3/2 or -3/2

OpenStudy (rulnick):

"By factoring": 4x^2-9 = 0 (2x-3)(2x+3) = 0 so 2x-3 = 0 gives x=3/2 and 2x+3 = 0 gives x=-3/2

OpenStudy (anonymous):

4x^2 = 9 (2x)^2 - 3^2 = 0 ( 2x -3) ( 2x + 3) = 0 x = + 3/2 , x = -3/2